### v6

@@ Abstract
Spatial attention has been widely used to improve the performance of convolutional neural networks. However, it has certain limitations. In this paper, we propose a new perspective on the effectiveness of spatial attention, which is that the spatial attention mechanism essentially solves the problem of convolutional kernel parameter sharing. However, the information contained in the attention map generated by spatial attention is not sufficient for large-size convolutional kernels. Therefore, we propose a novel attention mechanism called Receptive-Field Attention (RFA). Existing spatial attention, such as Convolutional Block Attention Module (CBAM) and Coordinated Attention (CA) focus only on spatial features, which does not fully address the problem of convolutional kernel parameter sharing. In contrast, RFA not only focuses on the receptive-field spatial feature but also provides effective attention weights for large-size convolutional kernels. The Receptive-Field Attention convolutional operation (RFAConv), developed by RFA, represents a new approach to replace the standard convolution operation. It offers nearly negligible increment of computational cost and parameters, while significantly improving network performance. We conducted a series of experiments on ImageNet-1k, COCO, and VOC datasets to demonstrate the superiority of our approach. Of particular importance, we believe that it is time to shift focus from spatial features to receptive-field spatial features for current spatial attention mechanisms. In this way, we can further improve network performance and achieve even better results. The code and pre-trained models for the relevant tasks can be found at https://github.com/Liuchen1997/RFAConv.
@@ 1 Introduction
Convolutional neural networks [1, 2] have dramatically reduced the computational overhead and complexity of models by using the convolutional operation with shared parameters. Driven by classical networks, such as LeNet [3], AlexNet [4], and VGG [5], convolutional neural networks have now established a complete system and formed advanced convolutional neural network models [6, 7, 8, 9, 10]. After carefully studying the convolutional operation, we gained inspiration. For classification, object detection, and semantic segmentation tasks, on one hand, the shape, size, color, and distribution of objects in different locations in the image are variable. However, during the convolutional operation, the convolutional kernel uses the same parameters in each receptive field to extract information, which does not consider the differential information from different locations. So the performance of the network is limited, as demonstrated by recent works [11, 12, 13]. On the other hand, the convolutional process does not take into account the significance of each feature, which further reduces the efficiency of the extraction features and ultimately restricts the performance of the model. In addition, the attention mechanism [14, 15, 16] enables the model to concentrate on significant features, which can enhance the benefits of feature extraction and the ability of convolutional neural networks to capture detailed feature information.
By examining the intrinsic limitations of convolutional operations and the properties of attention mechanisms, we assert that while the current spatial attention mechanism has fundamentally addressed the issue of parameter sharing in convolutional operations, but it remains restricted to the recognition of spatial features. The current spatial attention mechanism does not fully address the parameter sharing problem for larger convolutional kernels. Furthermore, they are unable to emphasize the significance of each feature in the receptive field, such as the existing Convolutional Block Attention Module (CBAM) [17] and Coordinate Attention (CA) [18]. Consequently, we propose a novel receptive-field attention (RFA) that comprehensively addresses the issue of parameter sharing for convolutional kernels and takes into account the significance of each feature in the receptive field. The RFA-designed convolutional operation (RFAConv) is a groundbreaking method that can replace standard convolutional operations in current neural networks. With only a few additional parameters and computational overhead, RFAConv enhances network performance. Numerous experiments conduct on ImagNet-1k [19], COCO [20], and VOC [21] have demonstrated the efficacy of RFAConv. As an attention-based convolutional operation, RFAConv outperforms CAMConv, CBAMConv, CAConv (constructed by CAM [17], CBAM, and CA, respectively), as well as the standard convolutional operation. Moreover, to address the issue of slow to extract receptive field features for current methods, we propose a lightweight operation. During the construction of RFAConv, we also design an upgraded version of CBAM and CA and conduct relevant experiments. We assert that spatial attention mechanisms should focus on receptive-field spatial features to further advance their development and enhance the advantages of convolutional neural networks.
@@ 2.1 Convolutional neural network architecture
The convolutional operation, which serves as a basic operation in convolutional neural networks, has led the development of many advanced network mode, such as vehicle detection [22], UAV images [23], medicine [24], etc. He et al. [25] suggested that as the depth of the network increases, the model becomes harder to train and may experience a degradation phenomenon. To address this issue, they proposed to use residual connections to revolutionize the design of network. Huang et al. [26] improved feature information by reusing features to address the issue of network gradient disappearance. After conducting a thorough study on convolutional operations, Dai et al. [27] claimed that the convolutional operation with a fixed sampling position can restrict the performance of the network to a certain extent. So, they proposed Deformable Conv, which alters the sampling positions of convolutional kernels by learning offsets. Building upon this approach, Deformable Conv V2 [28], and Deformable Conv V3 [29] were subsequently developed to further enhance the performance of convolutional networks. Zhang et al. [30] observed that group convolution could decrease the number of parameters and computational overhead of the model, however, insufficient interaction between information within the group can adversely affect the final network performance. Although the 1$\times$1 convolution can interact with information, it will bring more parameters and computational overhead, so they proposed the parameter-free Channel Shuffle operation to interact with the information between groups. Ma et al. [31] discovered that models with few parameters do not always result in faster inference times, and similarly, small computational effort does not guarantee quick performance. After careful study, they proposed the ShuffleNet V2. The YOLO [32] object detection network divided the input image into a grid to predict the location and class of objects. As research has progressed, eight versions of object detectors based on YOLO have been proposed, such as YOLOv5 [33], YOLOv7 [34], YOLOv8 [35], etc. While the previously-mentioned convolutional neural network architectures have achieved significant success, they do not directly address the problem of parameter sharing during the feature extraction process. Our work focuses on utilizing the attention mechanism to tackle the problem of convolutional parameter sharing from a fresh perspective.
@@ 2.2 Attention Mechanism
Attention mechanism, as a technique to improve the performance of network models, allows models to focus on key features. The theory of attention mechanism has now established a complete and mature system in the field of deep learning. Hu et al. [36] proposed a Squeeze-and-Excitation (SE) block to obtain the weights corresponding to each channel. This is achieved by compressing features to aggregate global channel information. Wang et al. [37] asserted that the correspondence between individual channels and weights is indirect when the SE interacts with information. Therefore, they designed the Efficient Channel Attention (ECA) by replacing the Fully Connected (FC) layer in the SE with a one-dimensional convolution of adaptive kernel size. Woo et al. [17] proposed the Convolutional Block Attention Module (CBAM), which combines channel attention and spatial attention. As a plug-and-play module, it can be embedded into convolutional neural networks to enhance network performance. Although SE and CBAM have allowed the network to achieve good performance, Hou et al. [18] found that the compressing feature in SE and CBAM lost too much information. Therefore, they proposed the lightweight coordinate attention (CA) to solve this problem. Fu et al. [38] designed a spatial attention module and a channel attention module to extend Fully Convolutional Networks (FCN) for modeling semantic interdependencies in the spatial and channel dimensions, respectively. Zhang et al. [39] generated feature maps at different scales on channels to build a more efficient channel attention mechanism. This paper introduces a new approach to address the issue of parameter sharing in standard convolutional operations. Our proposal involves combining attention mechanisms to create convolutional operations. Although existing attention mechanisms have demonstrated good performance, they do not specifically target the spatial features of receptive fields. To tackle this limitation, we developed RFAConv with non-shared parameters to improve performance of the network.
@@ 3 Methods
3.1 Reviewing Standard Convolutional Operation
Figure 1: It simply represents a 3$\times$3 convolution operation. The features are obtained by multiplying the convolution kernel with a receptive-field slider of the same size and then summing.
The standard convolutional operation serves as the fundamental building block for constructing convolutional neural networks. It utilizes sliding windows with shared parameters to extract feature information and overcome the inherent issues of neural networks constructed with fully connected layers, such as a large number of parameters and high computational overhead. Let X $\in$ $R^{C\times H\times W}$ denote the input feature maps, where C, H, and W represent the number of channels, height, and width of the feature map, respectively. To clearly demonstrate the feature extraction process through convolutional kernels, we use the example of C = 1. The convolutional operation to extract feature information from each receptive field slider can be expressed as follows:
$F_{1}=X_{11}\times K_{1}+X_{12}\times K_{2}+X_{13}\times K_{3}+...+X_{1S}\times K_{S}$ (1)
$F_{2}=X_{21}\times K_{1}+X_{22}\times K_{2}+X_{23}\times K_{3}+...+X_{2S}\times K_{S}$
$F_{N}=X_{N1}\times K_{1}+X_{N2}\times K_{2}+X_{N3}\times K_{3}+...+X_{NS}\times K_{S}$
Here, $F_{i}$ represents the value obtained by each convolutional slider after computation, $X_{i}$ represents the pixel value at the corresponding position within each slider, K represents the convolutional kernel, S denotes the number of parameters in the convolutional kernel, and N represents the total number of the receptive-field slider. As Shown in Figrue 1, K is a 3$\times$3 convolutional kernel and S is 9. Moreover, there are a total of 16 receptive-field sliders in the Figrue 1. For simplicity of representation, we just draw three receptive-field sliders. It can be seen that features at the same position within each slider share the same parameter $K_{i}$. Therefore, the standard convolutional operation fails to capture the difference in information brought by different positions, which ultimately limits the performance of the convolutional neural network to a certain extent.
3.2 Reviewing Spatial Attention
Currently, the spatial attention mechanism uses the attention map obtained through learning to highlight the importance of each feature. Similar to the previous section, C=1 is taken as an example. The spatial attention mechanism highlighting key features can be expressed simply as follows:
$F_{1}=X_{1}\times A_{1}$ (2)
$F_{2}=X_{2}\times A_{2}$
$F_{N}=X_{N}\times A_{N}$
Here, $F_{i}$ represents the value obtained after the weighting operation. $X_{i}$ and $A_{i}$ represent the values of the input feature map and the learned attention map at different positions, respectively, N is the product of the height and width of the input feature map, which represents the total number of pixel values. In general, the whole process can be simply represented in Fig. 2.
Figure 2: The original feature map highlights the key features by learned attention map. This process of highlighting is the Re-weight ($\times$) operation.
3.3 Spatial Attention and Standard Convolutional Operation
@@
As it is widely recognized, incorporating attention mechanisms into convolutional neural networks can enhance their performance. After a careful study of the standard convolution operation and the existing spatial attention mechanism, we argue that the spatial attention mechanism effectively overcomes the inherent limitation of convolutional neural networks, which is parameter sharing. Currently, the most commonly used kernel sizes in convolutional neural networks are 1 $\times$ 1 and 3 $\times$ 3. The convolutional operation for extracting features after introducing the spatial attention mechanism is either a 1 $\times$ 1 or a 3 $\times$ 3 convolutional operation. To intuitively show the process, the spatial attention mechanism is inserted into the front of the 1 $\times$ 1 convolutional operation. Through the attention map to weigh operation the input feature map (Re-weight ”$\times$”), finally, the slider feature information of the receptive-field is extracted through 1 $\times$ 1 convolution operation. The overall process can be simply represented as follows:
@@
Here, the convolution kernel K represents only one parameter value. If we take the value of $A_{i}\times K$ as a new convolution kernel parameter, the interesting thing is that the problem of parameter sharing in the extraction of features by 1$\times$1 convolution operations is solved. As shown in Fig. 3, it visualizes the combination of a 1$\times$1 convolutional kernel and a spatial attention mechanism. This combination essentially solves the problem of parameter sharing.
Figure 3: The convolutional kernel parameter $K_{i}$ obtained by multiplying the attentional weight $A_{i}$ with the convolutional kernel parameter K is different in each receptive-field slider, i.e., $Kernel1\neq Kernel2\neq Kernel3\neq...\neq KernelN$.
$F_{1}=X_{11}\times A_{11}\times K_{1}+X_{12}\times A_{12}\times K_{2}+X_{13}\times A_{13}\times K_{3}+...+X_{19}\times A_{19}\times K_{9}$ (4)
$F_{2}=X_{21}\times A_{21}\times K_{1}+X_{22}\times A_{22}\times K_{2}+X_{23}\times A_{23}\times K_{3}+...+X_{29}\times A_{29}\times K_{9}$
$F_{N}=X_{N1}\times A_{N1}\times K_{1}+X_{N2}\times A_{N2}\times K_{2}+X_{N3}\times A_{N3}\times K_{3}+...+X_{N9}\times A_{N9}\times K_{9}$
@@
Specifically, the output features are obtained by multiplying the receptive-field slider, the attention weights and the 1$\times$1 convolutional kernel parameters. In contrast to Fig. 1, if the product of the attention weights and the convolution kernel parameters is treated as a new convolution parameter, then the convolution parameter in each receptive field are different and not shared. However, when the spatial attention mechanism is inserted in front of the 3$\times$3 convolutional operation, it will be limited. As mentioned above, if we take the value of $A_{i}\times K$ as a new convolution kernel parameter, the Equation (4) completely solves the problem of parameter sharing for large-scale convolutional kernels. As shown in the Fig. 4, it represents the process in which the input features are weighted by spatial attention and then operated by a 3$\times$3 convolution. If the product of the attention weight corresponding to the receptive-field slider and the convolution kernel is taken as the convolution parameter, then the problem of convolution parameter sharing is solved. However, the most important point is that the convolutional kernel will share some of the features when it extracts features in each receptive field slider. In other words, there will be an overlap within each receptive-field slider, and when spatial attention is weighted, the size of the attention map is the same as the size of the input features, so the attention weights are shared in the sense field slider.
@@
After careful analysis will find that $A_{12}=A_{21},A_{13}=A_{22},A_{15}=A_{24},...$. In this case, the weights of the spatial attention map are shared across each sliding window. As a result, the spatial attention mechanism cannot effectively address the problem of parameter sharing for large-scale convolutional kernels, such as the 3$\times$3 convolution, because it does not consider the spatial features of the entire receptive field. Consequently, the effectiveness of the spatial attention mechanism is limited.
3.4 Innovating Spatial Attention and Standard Convolutional Operation
RFA is proposed to address the limitations of the existing spatial attention mechanism and provides an innovative solution to spatial processing. Drawing inspiration from RFA, a series of spatial attention mechanisms have been developed that can further enhance the performance of convolutional neural networks. RFA can be considered as a lightweight plug-and-play module, and the convolutional operation (RFAConv) designed by RFA can replace the standard convolution to improve the performance of convolutional neural networks.The spatial attention mechanism can solve the problem of convolutional parameter sharing by focusing on the receptive-field spatial feature, which is the direction for upgrading the existing spatial attention. The performance can be improved by combining the upgraded spatial features with the convolutional operation. So, we predict that the combination of spatial attention mechanisms and standard convolution operations will continue to evolve and lead to new breakthroughs in the future.
Receptive-Field Spatial Feature: In order to better understand the concept of receptive-field spatial feature, We will provide the relevant definition. The receptive-field spatial feature is specifically designed for convolutional kernels and is dynamically generated based on the kernel size. As shown in Fig. 5, the 3$\times$3 convolutional kernel is used as an example.
Figure 5: The receptive-field spatial features are obtained by transforming the spatial features.
In Figure 5, the ”Spatial Feature” refers to the original feature map. The ”Receptive-Field Spatial Feature” is the feature map transformed by spatial features, which is composed of non-overlapping sliding windows. Each 3$\times$3 size window in the receptive-field spatial feature represents a receptive-field slider.
Figure 6: In the figure, it shows in detail an example of extracting 3$\times$3 receptive-field spatial features by the Unfold method. For the convenience of display, we set the step size to 3.
Rceptive-Filed Attention Convolution (RFAConv): Regarding the receptive-field spatial feature, we propose receptive-field attention (RFA). This approach not only emphasizes the significance of different features within the receptive field slider, but also prioritizes the receptive-field spatial feature. Through this method, the problem of convolution kernel parameter sharing is completely solved. The receptive-field spatial features are dynamically generated depending on the size of the convolution kernel, therefore, The RFA is a fixed combination of convolution and cannot be separated from the help of convolution operations, which simultaneously rely on RFA to improve performance, so we propose receptive-field attention convolution (RFAConv). The overall structure of RFAConv with a 3$\times$3 size convolutional kernel is shown in Fig. 7.

### v7

@@ Abstract
In the realm of deep learning, spatial attention mechanisms have emerged as a vital method for enhancing the performance of convolutional neural networks. However, these mechanisms possess inherent limitations that cannot be overlooked. This work delves into the mechanism of spatial attention and reveals a new insight. It is that the mechanism essentially addresses the issue of convolutional parameter sharing. By addressing this issue, the convolutional kernel can efficiently extract features by employing varying weights at distinct locations. However, current spatial attention mechanisms focus on shallow attention to spatial features, which is insufficient to address the fundamental challenge of parameter sharing in convolutions involving larger kernels. In response to this challenge, we introduce a novel attention mechanism known as Receptive-Field Attention (RFA). Compared to existing spatial attention methods, RFA not only concentrates on the receptive-field spatial features but also offers effective attention weights for large convolutional kernels. Building upon the RFA concept, a Receptive-Field Attention Convolution (RFAConv) is proposed to supplant the conventional standard convolution. Notably, it offers nearly negligible increment of computational overhead and parameters, while significantly improving network performance. Furthermore, this work reveals that current spatial attention mechanisms require enhanced prioritization of receptive-field spatial features to optimize network performance. To validate the advantages of the proposed methods, we conduct many experiments across several authoritative datasets, including ImageNet, COCO, VOC, and Roboflow. The results demonstrate that the proposed methods bring about significant advancements in tasks, such as image classification, object detection, and semantic segmentation, surpassing convolutional operations constructed using current spatial attention mechanisms. Presently, the code and pre-trained models for the associated tasks have been made publicly available at https://github.com/Liuchen1997/RFAConv.
keywords:
attention mechanism, receptive-field attention, novel convolutional operation, classification, object detection, semantic segmentation.
{highlights}
In this work, the spatial attention mechanism is analyzed from a new perspective.
A novel attention mechanism is proposed, which solves the problem of convolution kernel parameter sharing.
RFAConv is proposed which can be considered as a high performance method to replace the $3~\times~3$ standard convolution.
RFCBAM and RFCA are integrated with standard convolution to develop non-shared parameter convolutional operations.
@@ 1 Introduction
Convolutional neural networks have dramatically reduced the computational overhead and complexity of models by using the convolutional operation with shared parameters. Currently, convolutional neural networks have now established a complete system and formed advanced convolutional neural networks, such as DCANet [1], HDTCNet[2] and FSENet [3].
However, standard convolution operations share parameters in each receptive-field slider, which limits network performance. Elsayed et al. [4] asserted that applying the same weights at all locations may be a less efficient use of computation than applying different weights at different locations. Jin et al. [5] argued that the standard convolution operation suffers from context-independence and that it should be fully aware of the specificity of each pixel. After carefully studying these works and standard convolutional operations, we gained inspiration. For the standard convolution operation, it uses the same parameters in each receptive-field to extract information, which does not take into account the difference information at different locations, as shown in Figure 1 (a). Therefore, the aim of this work is to consider the design of a non-shared parameter convolutional operation to perceive the difference in information at different locations, which is mainly achieved by assigning different convolutional parameters to the convolutional kernel at different locations, as depicted in Figure 1 (b) for clarity.
Figure 1: (a) Illustrates the process of extracting features from different locations using shared parameter convolution. (b) Illustrates the process of extracting features from different locations using non-shared parameter convolution. It is evident that information at different locations in the image varies, and using the same parameters to extract features does not fully consider the differentiation across different locations.
Building upon the aforementioned goals, we examine the inherent relationship between spatial attention mechanisms and conventional convolutions from a fresh standpoint. Upon analysis, we discover that the current spatial attention mechanism fundamentally address the issue of parameter sharing in convolutional operations, but it remains restricted to the recognition of spatial features. The current spatial attention mechanism does not fully address the parameter sharing problem for larger convolutional kernels. Furthermore, they are unable to emphasize the significance of each feature in the receptive-field, such as the existing Convolutional Block Attention Module (CBAM) [6] and Coordinate Attention (CA) [7].
In response to the aforementioned challenges, we introduce the innovative Receptive-Field Attention (RFA) mechanism, which not only effectively addresses the parameter sharing issue in convolution kernels but also meticulously considers the significance of each feature within the receptive-field, enabling precise feature capture and efficient utilization. Building upon the RFA concept, we introduce the Receptive-Field Attention Convolution (RFAConv) as a replacement for traditional standard convolution operations. With minimal additional parameters and computational overhead, RFAConv substantially enhances network performance, potentially supplanting traditional convolutional operations as a fundamental component in neural network architectures. Extensive experiments conducted on authoritative datasets like ImageNet, COCO, VOC, and Roboflow for tasks such as classification, object detection, and semantic segmentation, demonstrate that RFAConv yields significant improvements in image classification and object detection compared to convolutional operations utilizing existing spatial attention mechanisms. Moreover, we assert that the future advancement of spatial attention mechanisms should heavily emphasize receptive-field spatial features. By delving deep into this area’s potential, the strengths of convolutional neural networks will be further amplified. To substantiate this claim, the existing attention mechanism is optimized to seamlessly integrate receptive-field spatial features, effectively resolving the convolutional kernel parameter sharing issue. Therefore, building upon the foundational architectures of CBAM and CA, we propose two enhanced attention mechanisms RFCBAM and RFCA, and conduct comprehensive comparative experiments across various tasks, including classification, target detection, and semantic segmentation. The results strongly endorse our perspective, showcasing that attention mechanisms focusing on receptive-field spatial features can significantly enhance network performance once more.
@@ 2.1 Convolutional neural network architecture
After conducting a thorough study on convolutional operations, Dai et al. [8] claimed that the convolutional operation with a fixed sampling position can restrict the performance of the network to a certain extent. So, they proposed Deformable Conv, which alters the sampling positions of convolutional kernels by learning offsets. The convolutional operation, which serves as a basic operation in convolutional neural networks, has led the development of many advanced network mode. He et al. [9] suggested that as the depth of the network increases, the model becomes harder to train and may experience a degradation phenomenon. To address this issue, they proposed to use residual connections to revolutionize the design of network. Chen et al. [10] proposed the Coupled Global–Local (CGL) network for enhancing thee efficiency of object detection in large aerial images. It shares information from global branches and local branches, and can adaptively adjust the receptive-field for better feature extraction. Li et al. [11] devised the Large Selective Kernel Network (LSKNet), which can dynamically adjust its large spatial receptive-field to better model the ranging context of various objects in remote sensing scenarios. Ma et al. [12] discovered that models with few parameters do not always result in faster inference times, and similarly, small computational overhead does not guarantee quick performance. After careful study, they proposed the ShuffleNet V2. The YOLO object detection network divided the input image into a grid to predict the location and class of objects. As research has progressed, eight versions of object detectors based on YOLO have been proposed, such as YOLOv5 [13], YOLOv7 [14], YOLOv8 [15], etc. While the previously-mentioned convolutional neural network architectures have achieved significant success, they do not directly address the problem of parameter sharing during the feature extraction process. Our work focuses on utilizing the attention mechanism to tackle the problem of convolutional parameter sharing from a fresh perspective.
@@ 2.2 Attention Mechanism
Attention mechanism, as a technique to improve the performance of network models, allows models to focus on key features. The theory of attention mechanism has now established a complete and mature system in the field of deep learning. Wang et al. [16] asserted that the correspondence between individual channels and weights is indirect when the Squeeze-and-Excitation Block interacts with information. Therefore, they designed the Efficient Channel Attention (ECA) by replacing the Fully Connected (FC) layer in the SE with a one-dimensional convolution of adaptive kernel size. Woo et al. [6] proposed the Convolutional Block Attention Module (CBAM), which combines channel attention and spatial attention. As a plug-and-play module, it can be embedded into convolutional neural networks to enhance network performance. Although SE and CBAM have allowed the network to achieve good performance, Hou et al. [7] found that the compressing feature in SE and CBAM lost too much information. Therefore, they proposed the lightweight coordinate attention (CA) to solve this problem. Zhang et al. [17] generated feature maps at different scales on channels to build a more efficient channel attention mechanism. Misra et al. [18] proposed TripletAttention (TA), which de-transforms the dimensions of features and obtains three attention weights through three different branches. Finally, the branches are fused by weighting and averaging to obtain efficient feature. Yu et al. [19] argued that most existing approaches to attention mechanisms either ignore modelling attention in the channel and spatial dimensions, or introduce higher model complexity and greater computational burden. To alleviate this dilemma, they proposed a lightweight and efficient multidimensional collaborative attention (MCA), which simultaneously infers attention in the channel, height and width dimensions by using a three-branch architecture. Li et al. [20] introduced Hybrid Attention Module (HAM), which employs a channel attention module to simultaneously construct channel attention maps and produce channel-refined feature representations. Building upon these channel-wise attention weights, the spatial attention module subsequently partitions the feature maps along the channel dimension, generating complementary spatial attention descriptors through dual-path processing. These descriptors are then systematically integrated through attention-weighted feature modulation, enabling adaptive refinement of spatial feature representations.
This work introduces a new approach to address the issue of parameter sharing in standard convolutional operations. Our proposal involves combining attention mechanisms to create convolutional operations. Although existing attention mechanisms have demonstrated good performance, they do not specifically target the spatial features of receptive-fields. To tackle this limitation, we developed RFAConv with non-shared parameters to improve performance of the network.
@@ 3 Methods
Existing works predominantly concentrate on leveraging spatial attention to enhance network performance, exploring its contributions to feature extraction, information screening. While they indeed show the effectiveness of spatial attention to some degree, these analyses often approach spatial attention from a conventional viewpoint, overlooking the deeper relationship between its intrinsic mechanism and convolutional kernels. Therefore, this section analyses the intrinsic connection between spatial attention mechanisms and standard convolutional operations from a new perspective, and points out that existing spatial attention mechanisms do not adequately address the problem of parameter sharing for large convolutional kernels. To address this issue, we propose the concept of the receptive-field spatial feature, in which different attention weights can be learned for each receptive-filed slider. Based on this, we introduce a receptive-field attention convolution (RFAConv), which can be primarily segmented into Group Optimization and Receptive-Field Attention Process stages. Following these two processing steps, RFAConv effectively resolves the issue of parameter sharing to enhance network performance in classification, target detection, and semantic segmentation tasks.
3.1 Spatial Attention and Standard Convolutional Operations
@@
As is widely recognized, incorporating attention mechanisms into convolutional neural networks can enhance their performance. Regarding the effectiveness of spatial attention mechanisms, a new perspective is analyzed in this work. It effectively overcomes the inherent limitation of convolutional neural networks, which is parameter sharing. Currently, the most commonly used kernel sizes in convolutional neural networks are 1 $\times$ 1 and 3 $\times$ 3. The convolutional operation for extracting features after introducing the spatial attention mechanism is either a 1 $\times$ 1 or a 3 $\times$ 3 convolutional operation. To intuitively show the process, the spatial attention mechanism is inserted into the front of the 1 $\times$ 1 convolutional operation. The input feature map is weighted (Re-weight $\times"$) by the attention map. Finally, the slider feature information of the receptive-field is extracted through 1 $\times$ 1 convolution operation. The overall process can be simply represented as the Equation (LABEL:qe2). Due to the 1 $\times$ 1 convolution operation has only one convolution parameter, $K$ still represents only one parameter value, $A_{i}$ denotes a attention weight for the corresponding position, $F_{i}$ represents the value obtained by each convolutional slider after computation and N still represents the total number of the receptive-field slider. Unlike the 3 $\times$ 3 convolution, the 1 $\times$ 1 convolution has just one pixel value $X_{i}$ in each receptive-field slider.
@@
In the absence of spatial attention, each sliding window obtains $F_{i}$ simply by performing operations on $X_{i}$ with $K$. Therefore, analogous to standard convolution operations, if we take spatial attention weighting and convolution operations as an overall arithmetic step, the interesting thing is that the problem of parameter sharing in the extraction of features by 1 $\times$ 1 convolution operations is solved. In other words, the value of $A_{i}\times K$ is treated as a new convolution kernel parameter.
$F_{1}=X_{11}\times A_{11}\times K_{1}+X_{12}\times A_{12}\times K_{2}+X_{13}\times A_{13}\times K_{3}+.+X_{19}\times A_{19}\times K_{9}$ (2)
$F_{2}=X_{21}\times A_{21}\times K_{1}+X_{22}\times A_{22}\times K_{2}+X_{23}\times A_{23}\times K_{3}+.+X_{29}\times A_{29}\times K_{9}$
$F_{N}=X_{N1}\times A_{N1}\times K_{1}+X_{N2}\times A_{N2}\times K_{2}+X_{N3}\times A_{N3}\times K_{3}+.+X_{N9}\times A_{N9}\times K_{9}$
Here $X_{ij}$ and $A_{ij}$ represent the $j$-th feature value and $j$-th attentional weight in the $i$-th receptive-field slider. $K_{i}$ represents the $i$-th convolution parameter in the convolution kernel $K$. However, when the spatial attention mechanism is inserted in front of the 3 $\times$ 3 convolutional operation, it will be limited. As mentioned above, if we take the value of $A_{i}\times K$ as a new convolution kernel parameter, the Equation (2) completely solves the problem of parameter sharing for large-scale convolutional kernels. As shown in the Figure 2, it represents the process in which the input features are weighted by spatial attention and then operated by a 3 $\times$ 3 convolution. If the product of the attention weight corresponding to the receptive-field slider and the convolution kernel is taken as the convolution parameter, then the problem of convolution parameter sharing is solved.
@@
However, the most important point is that the convolutional kernel will share some of the features when it extracts features in each receptive-field slider. In other words, there will be an overlap within each receptive-field slider and when spatial attention is weighted, the size of the attention map is the same as the size of the input features, so the attention weights are shared in the receptive-field slider. Therefore, in Figure 2, there exists the case where $A_{12}=A_{21},A_{13}=A_{22},A_{15}=A_{24},~...~,A_{29}=A_{38}$. In this case, the weights of the spatial attention map are shared across each sliding window. As a result, the spatial attention mechanism cannot effectively address the problem of parameter sharing for large-scale convolutional kernels, such as the 3 $\times$ 3 convolution, because it does not consider the receptive-field spatial features. Consequently, the effectiveness of the spatial attention mechanism is limited.
3.2 Receptive-Filed Spatial Feature
As highlighted in the preceding subsection, current spatial attention mechanisms share a fraction of their attention weights across the receptive-field slider, resulting in a lack of effective addressing of the convolutional parameter sharing issue. To tackle this challenge, our approach centers on the receptive-field spatial features. It is specifically designed for convolutional kernels in this work and is dynamically generated based on the kernel size. As shown in Figure 3, the 3 $\times$ 3 convolutional kernel is used as an example.
Figure 3: The receptive-field spatial features are obtained by expanding the spatial features.
In Figure 3, the “Spatial Feature" refers to the original feature map. The “Receptive-Field Spatial Feature" is the feature map expanded by spatial features, which is composed of non-overlapping sliding windows. This is why the receptive-field spatial feature become three times wider and three times higher. In a subsequent section, we will describe in detail the method of expanding spatial features into receptive-field spatial features. Therefore, each 3 $\times$ 3 size window in the receptive-field spatial feature represents a receptive-field slider and is a non-overlapping sliding window. In this way, the weight of attention learned by focusing on the receptive-field spatial feature can not be shared in each receptive-field slider. Therefore after obtaining the attention weights and constructing a new operator with the standard convolution operation, then the problem of sharing the parameters of the standard convolution operation will be solved. Moreover, the spatial attention mechanism can solve the problem of convolutional parameter sharing by focusing on the receptive-field spatial feature, which is the direction for upgrading the existing spatial attention.
3.3 Receptive-Filed Attention Convolution
Regarding the receptive-field spatial feature, we propose Receptive-Filed Attention (RFA) to address the limitations of the existing spatial attention mechanism and provides an innovative solution to the standard convolutional operation. Through this method, each receptive-field slider will receive a different attentional weight. Thus its combination with the standard convolution operation will transform the convolution operation into a non-shared parameter operator to improve the performance of the convolutional neural network. Inspired by the above, we propose Receptive-Field Attention Convolution (RFAConv). The overall structure of RFAConv with a 3 $\times$ 3 size convolutional kernel is shown in Figure 4 and C = 3 is an example. Overall, the operational process of RFAConv can be divided into two steps, i.e., Grouping Optimization and Receptive-Field Attention Process.